OpenStudy (anonymous):

If I and H are midpoints of triangle EFG, find GE. A. 5 B. 10 C. 20 D.25

OpenStudy (anonymous):



OpenStudy (anonymous):

can you please help me understand this?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

which topic does this Q belong ?

OpenStudy (anonymous):

have no idea.

OpenStudy (anonymous):

do you understand it though?

hartnn (hartnn):

yeah, since I is mid-point of GE, IG=IE so, x+1 = 3x-2 can u find x from here ?

OpenStudy (anonymous):

my answer is either A B or C. x=3 over 2

OpenStudy (anonymous):



hartnn (hartnn):

corrrect! x= 3/2 so,GE = IG+IE = x+1+3x-2 = ... ?

OpenStudy (anonymous):

4x-1

OpenStudy (anonymous):

how do i get an answer from here?

hartnn (hartnn):

by putting x= 3/2

OpenStudy (anonymous):

i tried but couldnt get a right answer.. somethings off

hartnn (hartnn):

what did u try ? show steps ?

OpenStudy (anonymous):

i keep getting x=5/8

OpenStudy (anonymous):

4x-1=3/2 solve for x. x=5/8

hartnn (hartnn):

see, 4x-1 = 4 (3/2) -1 now 4/2 =2 so, 2*3-1 oh you solved for that, i actually told to put x=3/2 in 4x-1 :P

OpenStudy (anonymous):

then i tried 4x-1+3/2 but no luck either

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

my answers 5!

hartnn (hartnn):

mine too! :D

OpenStudy (anonymous):

thanks! :D

hartnn (hartnn):

most welcome ^_^

OpenStudy (anonymous):

i have a quick question. o: Why can't you go past 99 points on here?

hartnn (hartnn):

because there can be only one 100 here, and that position is taken by the respected satellite73 :)

OpenStudy (anonymous):

O: is that like the main boss? lol i never knew tht.

hartnn (hartnn):

he is like a math Guru he helped innumerable people here. he is here, from the beginning of OS

OpenStudy (anonymous):

oh snap. x) haha cool! and thanks again! :D

hartnn (hartnn):

welcome :)